OpenStudy (anonymous):

Three years ago, Andy invested $5,000 in an account that earns 5% interest compounded annually. The equation y = 5,000 (1.05)^t describes the balance in the account, where t is time in years. Andy made no additional deposits and no withdrawals. How much is in Andy's account at this time? a. $5, 250.26 b. $5, 470.19 c. $5, 750.00 d. $5, 788.13

OpenStudy (anonymous):

y=5,000*(1.05)^t y=5,000*(1.05)^3 y=5,000*(~1.158)-- that is approximately not a negative sign. y=5,788.125 so d.